### v1

@@ 5.2 Experiments with lognormal random fields
In this section, we consider a more realistic lognormal model for the uncertain coefficients, despite not yet having theoretical results as in the uniform case. The following lognormal model has been successfully employed in Bayesian inversion for gliomas in both murine models [30] and human patients [38],
@@
We empirically investigate the convergence rate using $R=16$ random shifts $\boldsymbol{\Delta}$} and $N=2^{m}$ lattice points with $m=[6,7,\dots 17]$. Each KL expansion is truncated at $512$ terms, for a stochastic dimension of $s=1024$. In the lognormal case, the shifted QMC points are ${\boldsymbol{y}}^{(i)}_{{\boldsymbol{\Delta}}}:=\boldsymbol{\Phi}^{-1}\left((\tfrac{i}{N}\,{\boldsymbol{z}}+{\boldsymbol{\Delta}})\bmod 1\right)$, where $\boldsymbol{\Phi}^{-1}(\cdot)$ denotes the inverse cumulative normal distribution applied componentwise to a vector. The same embedded lattice rule as in the uniform case is used to construct the QMC sequence. The empirical spectrum of the covariance operators $\mathcal{C}_{a}$ and $\mathcal{C}_{\kappa}$ and convergence results are shown in Figure 3. We observe that the QMC root-mean-square error converges at a faster rate of $\mathcal{O}(N^{-0.87})$ than the MC rate of $\mathcal{O}(N^{-1/2})$. The square-root of the eigenvalues decay like $k^{-1.17}$ and so we expect the QMC convergence rate to be $1/p-1/2\approx 1.17-1/2=0.67$, if the theory for uniform fields also holds for lognormal fields (as proven for some elliptic linear problems). Hence our results are very encouraging and prompt further analysis.
@@ Acknowledgments
The authors are grateful for insightful discussions with Albert Cohen, Wolfgang Dahmen, Mike Giles, Maria Pia Gualdani, Pierre L’Ecuyer, Dingcheng Luo, Fabio Nobile, Mark Peletier, William Porteous, and Christian Wieners, as well as other participants at the 2025 Oberwolfach Workshop on Computation and Learning in High Dimensions.

### v2

@@ 5.2 Experiments with lognormal random fields
In this section, we consider a more realistic lognormal model for the uncertain coefficients, despite not yet having theoretical results as in the uniform case. There is QMC theory to handle the lognormal case, see e.g., [20, 21], but it requires a new parametric regularity bound for the lognormal model of this problem which is anticipated to be harder and so is left for future work. The following lognormal model has been successfully employed in Bayesian inversion for gliomas in both murine models [33] and human patients [41],
@@
We empirically investigate the convergence rate using $R=16$ random shifts $\boldsymbol{\Delta}$} and $N=2^{m}$ lattice points with $m=[6,7,\dots 17]$. Each KL expansion is truncated at $512$ terms, for a stochastic dimension of $s=1024$. In the lognormal case, the shifted QMC points are ${\boldsymbol{y}}^{(i)}_{{\boldsymbol{\Delta}}}:=\boldsymbol{\Phi}^{-1}\left((\tfrac{i}{N}\,{\boldsymbol{z}}+{\boldsymbol{\Delta}})\bmod 1\right)$, where $\boldsymbol{\Phi}^{-1}(\cdot)$ denotes the inverse cumulative normal distribution applied componentwise to a vector. The same embedded lattice rule as in the uniform case is used to construct the QMC sequence. Computations are carried out in the same fashion as in Section 5.1. The reference calculation took 1,417 minutes (23.6 hours) and the QMC evaluations averaged 165 minutes (2.8 hours). We attribute the slight increase in computational cost to the larger stochastic dimension which resolves finer features as well as the non-uniform domain. The empirical spectrum of the covariance operators $\mathcal{C}_{a}$ and $\mathcal{C}_{\kappa}$ and convergence results are shown in Figure 3. We observe that the QMC root-mean-square error converges at a faster rate of $\mathcal{O}(N^{-0.86})$ than the MC rate of $\mathcal{O}(N^{-1/2})$. The square-root of the eigenvalues decay like $k^{-1.17}$ and so we expect the QMC convergence rate to be $1/p-1/2\approx 1.17-1/2=0.67$, if the theory for uniform fields also holds for lognormal fields (as proven for some elliptic linear problems). Hence our results are very encouraging and prompt further analysis.
@@ Acknowledgments
The authors give heartfelt thanks to both referees for their detailed suggestions to improve the paper, especially to one referee for explaining how to prove that the unique weak solution is always nonnegative. We are also grateful for insightful discussions with Albert Cohen, Wolfgang Dahmen, Mike Giles, Maria Pia Gualdani, Pierre L’Ecuyer, Dingcheng Luo, Fabio Nobile, Mark Peletier, William Porteous, and Christian Wieners, as well as other participants at the 2025 Oberwolfach Workshop on Computation and Learning in High Dimensions.